# Learning to Reconstruct Quirky Tracks

### v1

@@ Abstract
Analysis of data from particle physics experiments traditionally sacrifices some sensitivity to new particles for the sake of practical computability, effectively ignoring some potentially striking signatures. However, recent advances in ML-based tracking allow for new inroads into previously inaccessible territory, such as reconstruction of tracks which do not follow helical trajectories. This paper presents a demonstration of the capacity of ML-based tracking to reconstruct the oscillating trajectories of quirks. The technique used is not specific to quirks, and opens the door to a program of searching for many kinds of non-standard tracks.
@@ I Introduction
Hadronic collisions at the energy frontier have wide power to reveal new physics, such as new particles or interactions, whether expected or unexpected. The Large Hadron Collider (LHC), for example, has made important discoveries Aad et al. (2012); Chatrchyan et al. (2012) and placed tight constraints Aad et al. (2016); Cakir (2015); CMS (2016) on many models. In the coming years, the LHC will accumulate a dataset of unprecedented luminosity, with an expected ${\cal L}=300~{}{\rm fb}^{-1}$ ($3~{}{\rm ab}^{-1}$) with the completion of Run 3 (Run 4+). While the current physics program is broad, the technical challenges of analyzing the high-dimensional, high volume dataset have historically required imposing assumptions on the nature of the potential new physics, which translate to strategic sacrifices in sensitivity, especially to the unexpected. As a result, signatures of groundbreaking discoveries may have been overlooked by the traditional analysis pipelines.
@@
A crucial and challenging component of that pipeline is identifying tracks of charged particles in the inner, magnetized detector volume. Ionization leaves hits as particles pass through a tracking detector, but the task of assembling sets of hits to reconstruct tracks of particles is very challenging due to the enormous number of possible hit combinations. The task is greatly simplified, though still computationally costly, if one can assume that all tracks follow a helical trajectory and originate from within the beam pipe, as is expected of electrically charged particles created in an interaction at the beam crossing. This narrowed scope is sufficient for a vast number of expected signatures, but consequentially restricts the power to discover unexpected new physics. This is well understood, and significant effort has been made to relax the assumption that tracks must originate inside the beam pipe by developing dedicated software to reconstruct tracks with displaced verticesAad et al. (2023a); Lee et al. (2019); Alimena et al. (2020), as might be expected from decays of long-lived particles. Comparatively little effort Aad et al. (2023b); Abulencia et al. (2006); Knapen et al. (2017) has been put towards development of algorithms capable of identifying a potentially much broader class of tracks, those which are charged under other forces and do not follow helical paths in a solenoidal magnetic field. Such striking tracks may already exist in our data, but would be largely invisible to our current algorithms.
@@
An important challenge in reconstructing such non-helical tracks is that traditional tracking algorithms Strandlie and Fruhwirth (2010); Aaboud et al. (2017); Chatrchyan et al. (2014); Duda and Hart (1972) assume the specific parametric form of the desired track in a way that does not trivially generalize to new forms. Each new potential trajectory would therefore require its own bespoke tracking algorithm and incur its own computational expense in the analysis pipeline. However, recent innovations in machine-learning Bronstein et al. (2017) (ML) have allowed for the development of ML-based tracking algorithms which are architecturally agnostic to the parametric form of the track Ju et al. (2020a), where the target class of track is defined only by the sample used to train the algorithm. In principle, this allows for rapid development of tracking pipelines which are more flexible, replacing the complex work of dedicated algorithm development with the much simpler task of preparation of a new training sample. If successful, it lays the groundwork for a more ambitious task, preparation of a generalized training sample that produces a track finder capable of reconstructing a broad class of helical and non-helical trajectories, with true sensitivity to unexpected tracks.
@@
In this paper, we take the first step in that direction by demonstrating that ML-based tracking algorithms can be trained to identify non-helical tracks from a training sample, without the development of a dedicated new tracking algorithm. We train the graph-based machine learning algorithm Exa.TrkX Ju et al. (2021) to learn to reconstruct the strikingly non-helical oscillatory tracks produced by quirk pairs. Quirks Kang and Luty (2009) are charged under an unbroken confining QCD-like gauge symmetry but, in contrast with quarks, the quirk masses are large relative to the confinement scale, resulting in an oscillatory non-helical motion in a magnetic field for quirk pairs. These spectacular signatures are such a departure from the classical helical tracks, given the in-flight interactions between the quirks, that even a bespoke application of the traditional algorithms has been a long-standing challenge. Previous efforts have attempted to side-step this complexity by searching for straight tracks Abazov et al. (2010), momentum imbalance Farina and Low (2017), stopped out-of-time particles Evans and Luty (2019) or sets of hits along a plane Knapen et al. (2017), essentially ignoring the oscillatory nature of the tracks. This is reasonably effective for a limited quirk parameter range, but we show that ML-based tracking is more broadly efficient without requiring such quirk-specific simplifications. Our studies also demonstrate a capacity to generalize beyond examples in the training set. This gives hope that a similar approach could be used to construct a pipeline that can reconstruct tracks from a broader class of particles, unrestricted to helical tracks or quirks. This may pave the way for discoveries at the LHC that have evaded our detection. Complementary studies show sensitivity to quirks in the far-forward region with experiments such as FASER Feng et al. (2024); Li et al. (2021, 2023); Mammen Abraham et al. (2024).
@@
The paper is organized as follows. In Sec. II we cover the background of our study by reviewing the quirk model in a collider setting. In Sec. III, we discuss the equations of motion for the quirk pair and for the background particles. In Sec. IV, we discuss the Exa.TrkX GNN pipeline and how it can be applied to quirk track reconstruction. In Sec. V we present our our track reconstruction performance. In Sec. VI we conclude and discuss future directions.
@@
Figure 1: Example trajectories of quirk pairs in simulated events with a simplified detector geometry described in the text. The top row shows example trajectories for $\Lambda$ = 100 eV, and increases in mass from left to right, with $m_{Q}=500,1000,5000$ GeV. The center (bottom) row shows examples for the same mass range but for $\Lambda=500(1000)$ eV. Hits from SM background tracks are in grey, and from quirks are in blue. The true quirk trajectory is also shown, in red.
@@ II Theoretical Background
The unbroken ${\rm SU}(N)_{\rm IC}$ is leads to an infracolor gauge string between the pair-produced quirks, as with quarks. Unlike quarks however, as the quirks separate the breaking of the string is exponentially suppressed due to $m_{Q}\gg\Lambda$. The absence of string breaking results in oscillation of the pair with a characteristic length Knapen et al. (2017)
@@
in the center-of-mass frame, where $\gamma$ is the Lorentz boost of the quirks at production. When reporting the quirk track reconstruction performance we will use this formula for qualitative understanding. It is important to note that since we will be study pair produced quirks in association with a jet, actual oscillation lengths in the lab frame can be very different from Eq. 1 due to smaller-than-typical opening angles or different energies of each quirk.
In a collider setting, the oscillation length of produced quirks determines the potential reconstruction scenario. For $d_{cm}\ll 1$ cm, oscillation cannot be resolved by the detector, and the quirk pair is reconstructed as a single track. For oscillation scales much larger than the detector, $d_{cm}\gg 10$ m, the quirk pair does not complete an oscillation within the detector, and so the quirk pair appears as two standard tracks. In both cases, the tracks can be described using helical parameterization, traditional tracking algorithms are effective and existing searches are relevantKnapen et al. (2017); Evans and Luty (2019); Abazov et al. (2010); Farina and Low (2017).
However, when the oscillation length is on the order of the size of the detector, the quirk pair produces two tracks with a unique, non-helical motion. See Fig. 1 for visualizations of example trajectories. The stark deviations from a helix make traditional tracking algorithms ineffective for finding such complex tracks.
@@ III Modeling
Samples of simulated events are prepared and used for training of the tracking algorithm as well as evaluation of the performance. We study quirks which are charged under the Standard Model ${\rm U}(1)_{\rm EM}$ such that they can be pair produced through an $s$-channel Drell-Yan process. Quirk signal and SM background samples are generated with MadGraph5_aMC@NLO Alwall et al. (2014). Non-quirk particles are showered and hadronized with Pythia8 Bierlich et al. (2022) and a selection of $p_{\textrm{T}}>100~{}{\rm MeV}$ $|\eta|<2.5$ is applied on all charged particles, including quirks. The trajectories of particles and their interaction with the detector are then modeled with a custom package described below.
At leading order, the quirk pair has zero transverse momentum. For certain regions of the $(m_{Q},\Lambda)$ parameter space with oscillation lengths $\sim 1$ cm, the resultant motion is that the quirks drift in the ±$\pm~{}z$-direction Li et al. (2021) and only leave hits in the innermost layers of the detector. To generate visible tracks, we produce the quirk pair in association with a gluon from initial-state radiation, giving the quirk pair a net transverse momentum. We generate samples of simulated quirks in the mass range [100,5000] GeV. Fig. 2 shows the distributions of the simulated $\gamma$ factor of the quirk pair, $\gamma_{cm}$, quirk pair transverse momentum and opening angles in this range. These quantities broadly describe the oscillation that is calculated for quirk trajectories. In general, as the opening angle $\Delta\phi$ approaches $\pi$, and the quirk momenta approaches the kinematic limit, the resultant quirk motion will have the largest separation. An effect which is not shown in these figures is the case where $p_{Q1}\gg p_{Q2}$ (or vice versa). In this case $Q_{1}$ will drag $Q_{2}$ in the direction of $\vec{p_{1}}$ which will result in a complicated motion and exchange of momenta between each quirk.
In $pp$ collisions at $\sqrt{s}=13$ TeV, the production cross section for two quirks and a hard jet $\sigma_{QQj}\approx 1$ fb for $m_{Q}=500~{}{\rm GeV}$, which is sufficient to expect a significant number of quirks in the current LHC dataset. For the purposes of this study we will explore quirk masses up to 5 TeV; for the larger masses in this range, the cross section is too small for appreciable production at the LHC with current planned luminosities. Nevertheless, we explore these heavier cases to assess the strength of our reconstruction technique which may be useful at future colliders or for models with higher production cross sections. Larger cross sections can be achieved for quirks with lower $p_{\textrm{T}}$, with associated reconstruction challenges, or quirks charged under ${\rm SU}(3)_{\rm C}$.
@@
Instead of following a helical path, quirk pairs follow the motion of two objects subjected to a central force. Given that their angular momentum is approximately conserved, their trajectories are largely confined to a single planeKnapen et al. (2017). Solutions to the equations of motion assume an extra central force between the two quirks in addition to the Lorentz force that arises from the external electric and magnetic fields. Following Knapen et al. (2017), the system is modeled using a Nambu-Goto action with massive endpoints Luscher and Weisz (2002). This action leads to the equations of motion for quirks which includes the typical Lorenz contribution as well as a new contribution. In practice we solve the equation of motion in the quirk pair center of mass frame and boost back to the lab frame:
@@
$\frac{d(\gamma\vec{\beta}_{1})}{dt}=\frac{1}{m}\left[-\Lambda^{2}\left(\sqrt{1-v_{1\perp}^{2}}\,\vec{s}(\vec{x}_{1},\vec{x}_{2})+\frac{\vec{v}_{1\parallel}\cdot\vec{s}(\vec{x}_{1},\vec{x}_{2})\,\vec{v}_{1\perp}}{\sqrt{1-v_{1\perp}^{2}}}\right)+q\left(\vec{E}_{CM}+\frac{\vec{dx_{1}}}{dt}\times\vec{B}_{CM}\right)\right]$ (2)
$\frac{\vec{dx_{1}}}{dt}=\frac{\gamma\vec{\beta}_{1}}{\sqrt{1+(\gamma\vec{\beta}_{1})^{2}}}$ (3)
with similar equations for the second quirk, where $x_{i}$ is the location of quirk $i$, $\beta_{i}$ is the velocity, $m$ is the mass of each quirk, $\Lambda^{2}$ is the “string tension” between each quirk, $v_{\perp}$ and $v_{\parallel}$ are the perpendicular and parallel components of the quirk velocities respectively, $E_{CM}$ and $B_{CM}$ are the electric and magnetic fields in the detector in the quirk pair center of mass frame. Here, $\vec{s}(x_{1},x_{2})$ and $\vec{v}_{1\parallel}$ are defined as
$\vec{s}(\vec{x}_{1},\vec{x}_{2})=\frac{\vec{x}_{1}-\vec{x}_{2}}{||\vec{x}_{1}-\vec{x}_{2}||}$ (4)
$\vec{v}_{1\perp}=\frac{\vec{dx_{1}}}{dt}\vec{s}(\vec{x}_{1},\vec{x}_{2})^{2}\quad\quad\vec{v}_{1\parallel}=\frac{\vec{dx_{1}}}{dt}-\vec{v}_{1\perp}$ (5)
The equations of motion are numerically integrated to compute particle trajectories given the momenta.
@@
To assess the track reconstruction performance, we assume a few different simplified detector configurations for an LHC-style detector. The simplest is an 8-layer configuration with 4 tracking layers in the inner detector between $r=3.1$ cm and $r=12.25$ cm and 4 layers in the outer detector, between $r=29.9$ cm and $r=100$ cm. In a second configuration, we consider a configuration where there are 25 tracking layers equally spaced between $r=3.1$ cm and $r=53$ cm. We calculate particle trajectories and their intersection with a detector layer is recorded as a hit. No noise hits are included in these first studies, though hit-location uncertainties and pileup impact are assessed.
@@
Figure 1 shows selected examples of quirk trajectories for $\Lambda=100,500,1000$ eV (top row) and $m_{Q}$ = 500 GeV, 1 TeV, 5 TeV. The oscillation length depends strongly on the mass and $\Lambda$, as expected from Eq 1. As quirks become heavier, they naturally receive less kinetic energy which results in smaller oscillation lengths. For heavy quirks with large $\Lambda$, the oscillations can be quite rapid. Both very long and very short oscillation lengths are similar enough to traditional helical tracks that standard algorithms can have reasonable efficiency. The focus of this work is on the intermediate regime, where $d$ is approximately a few centimeters, making strikingly non-helical tracks.

### v2

@@ Abstract
Analysis of data from particle physics experiments traditionally sacrifices some sensitivity to new particles for the sake of practical computability, effectively ignoring some potentially striking signatures. However, recent advances in machine-learning (ML)-based tracking allow for new inroads into previously inaccessible territory, such as reconstruction of tracks that do not follow helical trajectories. This paper presents a demonstration of the capacity of ML-based tracking to reconstruct the oscillating trajectories of quirks, particles charged under an unbroken QCD-like gauge symmetry. The technique used is not specific to quirks, and opens the door to a program of searching for many kinds of non-standard tracks.
@@ I Introduction
Hadronic collisions at the energy frontier have wide power to reveal or constrain new physics Aad et al. (2012); Chatrchyan et al. (2012); Gehrmann and Malaescu (2022); Ferreira da Silva (2023); Aaij et al. (2015, 2014); Aad et al. (2010); Khachatryan et al. (2010), such as new particles or interactions, whether expected or unexpected. In the coming years, the Large Hadron Collider (LHC) will accumulate a dataset of unprecedented luminosity, with an expected ${\cal L}=300~{\rm fb}^{-1}$ ($3~{\rm ab}^{-1}$) with the completion of Run 3 (Run 4+). While the current physics program is broad, the technical challenges of analyzing the high-dimensional, high volume dataset have historically required imposing assumptions on the nature of the potential new physics, which translate to strategic sacrifices in sensitivity, especially to the unexpected. As a result, signatures of groundbreaking discoveries may be overlooked by the traditional analysis pipelines.
@@
A crucial and challenging component of that pipeline is identifying tracks of charged particles in the inner, magnetized detector volume. Ionization leaves hits as particles pass through a tracking detector, but the task of assembling sets of hits to reconstruct tracks of particles is a challenging combinatorial problem. The task is made tractable, though still computationally costly, if one can assume that all tracks follow a helical trajectory and originate from within the beam pipe, as is expected of electrically charged particles created in an interaction at the beam crossing. This narrowed scope is sufficient for many expected signatures, but consequentially restricts the power to discover unexpected new physics. This is well understood, and significant effort has been made to relax the assumption that tracks must originate inside the beam pipe by developing dedicated software to reconstruct tracks with displaced vertices (recent results include Refs.Hayrapetyan et al. (2024a); Aad et al. (2023a); Hayrapetyan et al. (2024b); Lee et al. (2019); Alimena et al. (2020); Hayrapetyan et al. (2025a)), as might be expected from decays of long-lived particles. Comparatively little effort Aad et al. (2023b); Hayrapetyan et al. (2025b); Abulencia et al. (2006); Knapen et al. (2017); Hayrapetyan et al. (2025b) has been put towards development of algorithms capable of identifying a potentially much broader class of tracks, those which are charged under other forces and do not follow helical paths in a solenoidal magnetic field. Such striking tracks may already exist in our data, but would be largely invisible to our current algorithms.
@@
An important challenge in reconstructing such non-helical tracks is that traditional tracking algorithms Strandlie and Fruhwirth (2010); Aaboud et al. (2017); Chatrchyan et al. (2014); Duda and Hart (1972) assume the specific parametric form of the desired track in a way that does not trivially generalize to new forms. Each new potential trajectory would therefore require its own bespoke tracking algorithm and incur its own computational expense in the analysis pipeline. However, recent innovations in machine learning (ML) such as geometric deep learning Bronstein et al. (2017) have allowed for the development of ML-based tracking algorithms which are architecturally agnostic to the parametric form of the track Ju et al. (2020a), where the target class of track is defined only by the sample used to train the algorithm. In principle, this allows for rapid development of tracking pipelines which are more flexible, replacing the complex work of dedicated algorithm development with the much simpler task of preparation of a new training sample. If successful, it lays the groundwork for a more ambitious task, preparation of a generalized training sample that produces a track finder capable of reconstructing a broad class of helical and non-helical trajectories, with true sensitivity to unexpected tracks.
@@
In this paper, we take the first step in that direction by demonstrating that ML-based tracking algorithms can be trained to identify non-helical tracks from a training sample, without the development of a dedicated new tracking algorithm. We train the graph-neural-network (GNN) machine learning algorithm Exa.TrkX Ju et al. (2021) to learn to reconstruct the strikingly non-helical oscillatory tracks produced by quirk pairs. Quirks Kang and Luty (2009) are charged under an unbroken confining QCD-like gauge symmetry but, in contrast with quarks, the quirk masses are large relative to the confinement scale, resulting in an oscillatory non-helical motion by quirk pairs in a magnetic field. These spectacular signatures are such a departure from the classical helical tracks, given the in-flight interactions between the quirks, that even a bespoke application of the traditional algorithms has been a long-standing challenge. Previous efforts have attempted to side-step this complexity by searching for straight tracks Abazov et al. (2010), momentum imbalance Farina and Low (2017), stopped out-of-time particles Evans and Luty (2019) or sets of hits along a plane Knapen et al. (2017), essentially ignoring the oscillatory nature of the tracks. This is reasonably effective for a limited quirk parameter range, but we show that ML-based tracking is more broadly efficient without requiring such quirk-specific simplifications. Our studies also demonstrate a capacity to generalize beyond examples in the training set. This gives hope that a similar approach could be used to construct a pipeline that can reconstruct tracks from a broader class of particles, unrestricted to helical tracks or quirks. This may pave the way for discoveries at the LHC that have evaded our detection. Complementary studies show sensitivity to quirks in the far-forward region with experiments such as FASER Feng et al. (2024); Li et al. (2021, 2023); Mammen Abraham et al. (2024).
@@
The paper is organized as follows. In Sec. II we cover the background of our study by reviewing the quirk model in a collider setting. In Sec. III, we discuss the equations of motion for the quirk pair and for the background particles. In Sec. IV, we discuss the Exa.TrkX pipeline and how it can be applied to quirk track reconstruction. In Sec. V we present our our track reconstruction performance. In Sec. VI we discuss the sensitivity of our reconstruction techniques to quirks. In Sec. VII we conclude and discuss future directions.
@@
Figure 1: Example trajectories of quirk pairs in simulated events with a simplified detector geometry described in the text. The top row shows example trajectories for $\Lambda$ = 100 eV, and increases in mass from left to right, with $m_{Q}=500,1000,5000$ GeV, all with small oscillation distances. The center (bottom) row shows examples for the same mass range but for $\Lambda=500(1000)$ eV, both with visible oscillations. Hits from SM background tracks (see Sec V.4 for details) are in grey, and from quirks are in blue. The true quirk trajectory is also shown, in red.
@@ II Theoretical Background
The unbroken ${\rm SU}(N)_{\rm IC}$ leads to an infracolor gauge string between the pair-produced quirks, as with quarks. Unlike quarks however, as the quirks separate the breaking of the string is exponentially suppressed due to $m_{Q}\gg\Lambda$. The absence of string breaking results in oscillation of the pair with a characteristic length Knapen et al. (2017)
@@
in the center-of-mass frame, where $\gamma$ is the Lorentz boost of the quirks at production. When reporting the quirk track reconstruction performance we will use this formula for qualitative understanding. It is important to note that since we study pair produced quirks in association with a jet, actual oscillation lengths in the lab frame can be very different from Eq. 1 due to smaller-than-typical opening angles or different energies of each quirk.
In a collider setting, the oscillation length of produced quirks determines the potential reconstruction scenario. For $d_{cm}\lesssim 1$ mm, oscillation cannot be resolved by the detector, and the quirk pair is reconstructed as a single track. For oscillation scales much larger than the detector, $d_{cm}\gtrsim 1$ m, the quirk pair does not complete an oscillation within the detector, and so the quirk pair appears as two standard tracks. In both cases, the tracks can be described using helical parameterization, traditional tracking algorithms are effective and existing searches are relevantKnapen et al. (2017); Evans and Luty (2019); Abazov et al. (2010); Farina and Low (2017).
However, when the oscillation length is on the order of the size of the detector, the quirk pair produces two tracks with a unique, non-helical motion. See Fig. 1 for visualizations of example trajectories, whose features are explained in greater detail in the following section. The stark deviations from a helix make traditional tracking algorithms ineffective for finding such complex tracks.
@@ III Modeling
Samples of simulated events are prepared and used for training of the tracking algorithm as well as evaluation of the performance. We study quirks charged under the Standard Model ${\rm U}(1)_{\rm EM}$ (with charge $\pm 1$) such that they can be pair produced through, e.g., an $s$-channel Drell-Yan process. Samples of simulated quirk signal and SM background are generated in $pp$ collisions at $\sqrt{s}=13$ TeV with MadGraph5_aMC Alwall et al. (2014). Non-quirk particles are showered and hadronized with Pythia8 Bierlich et al. (2022) and a selection of $p_{\textrm{T}}>100~{\rm MeV}$ and $|\eta|<2.5$ is applied on all charged particles, including quirks. The trajectories of particles and their interaction with the detector are then modeled with a custom package described below.
The leading order Drell-Yan production mechanism results in a quirk pair with zero transverse momentum, but each quirk can have nonzero transverse momentum. For certain regions of the $(m_{Q},\Lambda)$ parameter space with oscillation lengths $\sim 1$ cm, the resultant motion is that the quirks travel in the beam direction Li et al. (2021) and only leave hits in the innermost layers of the detector. To generate visible tracks, we produce the quirk pair in association with a jet, giving the quirk pair a net transverse momentum. We generate samples of simulated quirks in the mass range $m_{Q}\in$[100,5000] GeV. Fig. 2 shows the distributions of the simulated $\gamma$ factor of the quirk pair, $\gamma_{cm}$, quirk pair transverse momentum and opening angles in this range; here, the opening angle refers to the initial angle between the quirks within the plane of their production. These quantities broadly describe the oscillation calculated for quirk trajectories. In general, as the opening angle, $\Delta\phi$, approaches $\pi$, and for quirks produced with large momenta, the resultant quirk motion will have the largest separation. An additional feature of the initial quirk production which has important consequences for the quirk trajectory is the case where $p_{Q1}\gg p_{Q2}$ (or vice versa). In this case $Q_{1}$ will drag $Q_{2}$ in the direction of $\vec{p_{1}}$ which will result in a complicated motion and exchange of momenta between each quirk.
In $pp$ collisions at $\sqrt{s}=13$ TeV, the production cross section for two quirks and a jet $\sigma_{QQj}\approx 1$ fb for $m_{Q}=500~{\rm GeV}$, which is sufficient to expect a significant number of quirks in the current LHC dataset. For the purposes of this study we will explore quirk masses up to 5 TeV; for the larger masses in this range, the cross section is too small for appreciable production at the LHC with current planned luminosities. Nevertheless, we explore these heavier cases to assess the strength of our reconstruction technique which may be useful at future colliders or for models with higher production cross sections. Larger cross sections can be achieved for quirks with lower $p_{\textrm{T}}$, with associated reconstruction challenges, or quirks charged under ${\rm SU}(3)_{\rm C}$.
@@
Instead of following a helical path, quirk pairs follow the motion of two objects subjected to a central force. Given that their angular momentum is approximately conserved, their trajectories are largely confined to a single planeKnapen et al. (2017). Solutions to the equations of motion assume an extra central force between the two quirks in addition to the Lorentz force that arises from the external electric and magnetic fields. Following Ref. Knapen et al. (2017), the system is modeled using a Nambu-Goto action with massive endpoints Luscher and Weisz (2002), leading to the equations of motion for a quirk given by the force:
@@
$\frac{d(m\gamma\vec{v})}{dt}=-\Lambda^{2}\left(\frac{\sqrt{1-v_{\perp}^{2}}}{v_{\parallel}}\,\vec{v}_{\parallel}+\frac{v_{\parallel}}{\sqrt{1-v_{\perp}^{2}}}\vec{v}_{\perp}\right)+\vec{F}_{\rm ext},$ (2)
with a similar equation for the second quirk. Here, $m$ is the mass of the quirk, $\vec{v}_{\perp}$ and $\vec{v}_{\parallel}$ are the perpendicular and parallel velocities of the quirk with respect to the infracolor string, and $F_{\rm ext}$ is the external force acting on the quirks, which for us will only include the affect of the external magnetic field on the (anti-)quirk with charge $q=\pm 1$.
@@
To assess the track reconstruction performance, we assume two simplified detector configurations for an LHC-style detector. The simplest is an 8-layer Knapen et al. (2017) configuration with 4 tracking layers in the inner detector between $r=3.1$ cm and $r=12.25$ cm covering $|z|<40~{\rm cm}$ and 4 layers in the outer detector, between $r=29.9$ cm and $r=100$ cm, covering $|z|<75~{\rm cm}$. In a second configuration, we consider a configuration where there are 25 tracking layers between $r=3.1$ cm and $r=53$ cm with the same $z$-range as the 8 -layer detector. We calculate particle trajectories and their intersection with a detector layer is recorded as a hit. No noise hits are included in these first studies, though hit-location uncertainties are assessed in Sec. V.8.
@@
Figure 1 shows selected examples of quirk trajectories for $\Lambda=100,500,1000$ eV (top row) and $m_{Q}$ = 500 GeV, 1 TeV, 5 TeV. The oscillation length depends strongly on the mass and $\Lambda$, as expected from Eq 1. As quirks become heavier, they naturally receive less kinetic energy which results in smaller oscillation lengths. For heavy quirks with large $\Lambda$, the oscillations can be quite rapid. Both very long and very short oscillation lengths are similar enough to traditional helical tracks that standard algorithms can have reasonable efficiency. The focus of this work is on the intermediate regime, where the quirk separation is approximately a few centimeters, making strikingly non-helical tracks.